# Tailoring transport in quantum spin chains via disorder and collisions

### v2

@@ 1 Introduction
The paper is organized as follows. In Section 2 we present the model of our disordered spin-chain interacting with the environment via stochastic collision models and introduce the figures of merit adopted to study transport properties. In Section 3, we present the results of our study comparing the cases of one, two and multiple injected excitations also highlighting the entanglement properties in those regimes. Finally, in Section 4 we present our conclusions and discuss future perpectives in the light of technological progress in quantum science and biology.
2 Model and methods
@@
$H_{XXZ}=J\sum_{i=1}^{N-1}\left[\sigma^{x}_{i}\sigma^{x}_{i+1}+\sigma^{y}_{i}\sigma^{y}_{i+1}+\Delta\sigma^{z}_{i}\sigma^{z}_{i+1}\right]+h\sum_{i=1}^{N}r_{v}(1,i)\sigma^{z}_{i},$ (1)
and subjected to dissipation through noise in the form of stochastic collision models, as depicted in Sec. 2.1, with open boundary conditions (OBC). In Eq. 1, $\sigma^{x/y/z}_{i}$ are the Pauli operators acting on spin located at site $i$; $J$ is the spin exchange rate, that can also be seen as the tunneling amplitude for excitation quasiparticles throughout the system; $\Delta$ governs the anisotropy in the form of a spin-spin interaction along the z-axis. In our analysis, we consider the regime with $J=1(>0)$ and then scale every other energy to $J$. As a result, the ground state with disorder $h=0$ is a ferromagnet for $\Delta>1$, becoming a critical paramagnet for $\Delta\leq 1$.
@@ 2.1 Noise modeling
with $U_{\textrm{coll}}=\exp{-i(\pi/2)\sigma_{a}^{x}\otimes\sigma^{z}_{i}}$ ) representing the collision event and $t_{i}$ being the time before the given collision. For a detailed description, we refer to our previous work [16].
@@ 2.2 Characterizing transport
3 Results
3.1 One excitation case
@@
Figure 2: (De)localization behaviour: one excitation case. Inverse Participation Ratio (IPR) for a spin chain with $N=41$ sites. (a) IPR at long time ($tJ=30$), for $h=10$ and different shape parameters $\nu$ and collision rates $r_{c}$. (b) IPR vs time for large disorder range $h=10$ for different collision rates $r_{c}$. (c) IPR vs time $tJ$ for low $r_{c}=0.1$ and different disorder range $h$. Notice the emergence of plateaus of temporary stabilization of localized behavior. For (b) and (c) the shape parameter is $\nu=100$ in the time-homogeneous collision regime. The data refer to the case of no anisotropy ($\Delta=0$ in Eq. 1) since they are seen to be not significantly altered by finite $\Delta\neq 0$. The simulations were performed with $M=500$ trajectories and $dt=0.02$.
@@ 3.2 Multiple excitations
3.2.1 Effect of anisotropy for two excitations
@@
Figure 3: Effect of anisotropy: the case of two excitations. Inverse Ergodicity ratio (IER) with associated magnetization, plateau length and delocalization time $\tau$ for a spin chain of $N=20$ sites and final time of simulations $tJ=30$. (a) IER vs time $tJ$ for different collision rates $r_{c}$ with anisotropy strength $\Delta=2.5$. Inset: magnetization map over sites $i$ and time $tJ$, showing the spreading of the two excitations over time. The two neighbouring excitations are initially separated by two spins with very homogeneous collisions over time ($\nu=100$) and high disorder range with $h=10$. The IER has the same trend of the IPR and we focus here on the collision rate regime $0\leq r_{c}\leq 1$, proved to be more interesting due to the formation of plateaus. For lower collision rates ($r_{c}\ll 1$) the systems remains localized (IER closer to 1), while increasing $r_{c}$ leads to a faster delocalization (IER decaying to its limit). (b) IER vs $tJr_{c}$ varying the collision rate parameter $0\leq r_{c}\leq 1$, at fixed high time-homogeneity ($\nu=100$), high disorder strength ($h=10$), and in the presence of anisotropy, considering two excitations initially separated by two spins.
We see that all curves collapse in the region where plateaus are present since they are naturally shifted with respect to each other. (c) 3D plot of plateau length $L$ in terms of tJ vs disorder range $h\in[0,10]$ and collision rate $r_{c}\in[0,1]$. (d) 3D plot of the complete delocalization time $\tau$ vs $h\in[0,10]$ and $r_{c}\in[0,10]$. The plane $r_{c}=h$ separates high $\tau$ regions from the small $\tau$ ones. In (c) and (d) different colors and shapes for the symbols represent different levels of collision time-homogeneity (colors) and anisotropy (shapes) as in the legend. Also, we show data from two excitations initially injected both next to each other and separated by two spins. The simulations were performed with $M=250$ trajectories and $dt=0.02$.
@@ 3.2.2 The full picture
To this aim, we shrink our spin chain to $N=8$ and inject $N_{exc}=4$ into the system, starting all next to each other to take into account the case where anisotropy plays a role. As we learned from Sec. 3.2.1 indeed, we restrict the parameter range to the most significant behavior. To sum up, this is for space-homogeneous (e.g. $\nu=100$) collisions happening at low rates ($0.1\leq r_{c}\leq 5)$, high disorder range bandwidth (e.g. $h=10$) and anisotropy as large as $\Delta=2.5$.
@@
3.3 Entanglement Entropy
To complement the studies based on local observables, in this section we analyze the behavior of entanglement as a measure for thermalization dynamics in our spin chain. We focus on the case of two excitations injected into the system next to each other, that is the most interesting situation found in previous sections, where disorder, noise and coupling between spins significantly affected the delocalization behavior. We limit to two excitations for computational ease, having now clear what does change from two to more. Finally, we again zoom the parameters in the interesting range emerged in Sec. 3.2.1, that allows the emergence of plateaus and non-continuous temporal profiles of transport: this is low collision rates, e.g. with $r_{c}=0.1$, large disorder range e.g. with $h=10$ and important anisotropy e.g. with $\Delta=2.5$. Finally, we now want to check the dependence on the noise space-homogeneity by letting the shape parameter $\nu$ vary from absence of noise ($\nu=0$), heterogeneous collisions over time ($\nu\leq 1$) and time-homogeneous collisions ($\nu=100$).
@@ 4 Discussion
From the simplest one excitation case, see Sec. 3.1, we found that disorder tends to suppress transport through interference, which competes with collisional noise while the latter introduces dephasing, thus facilitating delocalization, instead. One first relevant message emerging from our study is that among these competing mechanisms, low collision rates ($r_{c}\ll 1$) favor the occurrence of localized regions in the form of Inverse Participation Ratio (IPR) plateaus in time windows where collisions are absent. Interestingly, the length and rate of these plateaus depend directly on noise parameters with a universal scaling, allowing to systematically tailor sequences of transport and localization that can be applied for example to stroboscopic protocols in quantum technologies. At higher collision rates ($r_{c}\geq 1$) instead, frequent collisional events completely disrupt localization, driving the system toward a diffusive regime characterized by smooth evolution of IPR over time, even for moderate levels of disorder range.
From the case of several excitations, see Sec. 3.2 and Appendix B, we learn that additional features emerge due to interactions among excitations. In particular, more localized regions in the form of plateaus appear even for the tiniest levels of disorder range.
@@ Appendix A. Imbalance
Appendix B. Two excitations next to each other

### v3

@@ I Introduction
The paper is organized as follows. In Section II we present the model of our disordered spin-chain interacting with the environment via stochastic collision models and introduce the figures of merit adopted to study transport properties. In Section III, we present the results of our study comparing the cases of one, two and multiple injected excitations also highlighting the entanglement properties in those regimes. Finally, in Section IV we present our conclusions and discuss future perpectives in the light of technological progress in quantum science and biology.
II Model and methods
@@
$H_{XXZ}=J\sum_{i=1}^{N-1}\left[\sigma^{x}_{i}\sigma^{x}_{i+1}+\sigma^{y}_{i}\sigma^{y}_{i+1}+\Delta\sigma^{z}_{i}\sigma^{z}_{i+1}\right]$
$+h\sum_{i=1}^{N}r_{v}(1,i)\sigma^{z}_{i},$ (1)
and subjected to dissipation through noise in the form of stochastic collision models, as depicted in Sec. II.1, with open boundary conditions (OBC). In Eq. 1, $\sigma^{x/y/z}_{i}$ are the Pauli operators acting on spin located at site $i$; $J$ is the spin exchange rate, that can also be seen as the tunneling amplitude for excitation quasiparticles throughout the system; $\Delta$ governs the anisotropy in the form of a spin-spin interaction along the z-axis. In our analysis, we consider the regime with $J=1(>0)$ and then scale every other energy to $J$. As a result, the ground state with disorder $h=0$ is a ferromagnet for $\Delta>1$, becoming a critical paramagnet for $\Delta\leq 1$.
@@ II.1 Noise modeling
with $U_{\textrm{coll}}=\exp{-i(\pi/2)\sigma_{a}^{x}\otimes\sigma^{z}_{i}}$ representing the collision event and $t_{i}$ being the time before the given collision. For a detailed description, we refer to our previous work [16].
@@ II.2 Characterizing transport
III Results
III.1 One excitation case
@@
Figure 2: (De)localization behaviour: one excitation case. Inverse Participation Ratio (IPR) for a spin chain with $N=41$ sites. (a) IPR at long time ($tJ=30$), for $h=10$ and different shape parameters $\nu$ and collision rates $r_{c}$. (b) IPR vs time for large disorder range $h=10$ for different collision rates $r_{c}$. Curves with progressively increasing tickness and color scale refer to increasing values of $r_{c}=0,0.1,0.5,1,5,50,100$ (c) IPR vs time $tJ$ for low $r_{c}=0.1$ and different disorder range $h$. Notice the emergence of plateaus of temporary stabilization of localized behavior. For (b) and (c) the shape parameter is $\nu=100$ in the time-homogeneous collision regime. The data refer to the case of no anisotropy ($\Delta=0$ in Eq. 1) since they are seen to be not significantly altered by finite $\Delta\neq 0$. The simulations were performed with $M=500$ trajectories and $dt=0.02$.
@@ III.2 Multiple excitations
III.2.1 Effect of anisotropy for two excitations
@@
Figure 3: Effect of anisotropy: the case of two excitations. Inverse Ergodicity ratio (IER) with associated magnetization, plateau length and delocalization time $\tau$ for a spin chain of $N=20$ sites and final time of simulations $tJ=30$. (a) IER vs time $tJ$ for different collision rates $r_{c}$ with anisotropy strength $\Delta=2.5$. Inset: magnetization map over sites $i$ and time $tJ$, showing the spreading of the two excitations over time. The two neighbouring excitations are initially separated by two spins with very homogeneous collisions over time ($\nu=100$) and high disorder range with $h=10$. The IER has the same trend of the IPR and we focus here on the collision rate regime $0\leq r_{c}\leq 1$, proved to be more interesting due to the formation of plateaus. For lower collision rates ($r_{c}\ll 1$) the systems remains localized (IER closer to 1), while increasing $r_{c}$ leads to a faster delocalization (IER decaying to its limit). (b) IER vs $tJr_{c}$ varying the collision rate parameter $0\leq r_{c}\leq 1$, at fixed high time-homogeneity ($\nu=100$), high disorder strength ($h=10$), and in the presence of anisotropy, considering two excitations initially separated by two spins. Curves with progressively increasing thickness refer to values of $r_{c}$ increasing from 0.20 to 1.00 by 0.05 each. We see that all curves collapse in the region where plateaus are present since they are naturally shifted with respect to each other. (c) 3D plot of plateau length $L$ in terms of tJ vs disorder range $h\in[0,10]$ and collision rate $r_{c}\in[0,1]$. (d) 3D plot of the complete delocalization time $\tau$ vs $h\in[0,10]$ and $r_{c}\in[0,10]$. The plane $r_{c}=h$ separates high $\tau$ regions from the small $\tau$ ones. In (c) and (d) different colors and shapes for the symbols represent different levels of collision time-homogeneity (colors) and anisotropy (shapes) as in the legend. Also, we show data from two excitations initially injected both next to each other and separated by two spins. The simulations were performed with $M=250$ trajectories and $dt=0.02$.
@@ III.2.2 The full picture
To this aim, we shrink our spin chain to $N=8$ and inject $N_{exc}=4$ into the system, starting all next to each other to take into account the case where anisotropy plays a role. As we learned from Sec. III.2.1 indeed, we restrict the parameter range to the most significant behavior. To sum up, this is for space-homogeneous (e.g. $\nu=100$) collisions happening at low rates ($0.1\leq r_{c}\leq 5)$, high disorder range bandwidth (e.g. $h=10$) and anisotropy as large as $\Delta=2.5$.
@@
III.3 Entanglement Entropy
To complement the studies based on local observables, in this section we analyze the behavior of entanglement as a measure for thermalization dynamics in our spin chain. We focus on the case of two excitations injected into the system next to each other, that is the most interesting situation found in previous sections, where disorder, noise and coupling between spins significantly affected the delocalization behavior. We limit to two excitations for computational ease, having now clear what does change from two to more. Finally, we again zoom the parameters in the interesting range emerged in Sec. III.2.1, that allows the emergence of plateaus and non-continuous temporal profiles of transport: this is low collision rates, e.g. with $r_{c}=0.1$, large disorder range e.g. with $h=10$ and important anisotropy e.g. with $\Delta=2.5$. Finally, we now want to check the dependence on the noise space-homogeneity by letting the shape parameter $\nu$ vary from absence of noise ($\nu=0$), heterogeneous collisions over time ($\nu\leq 1$) and time-homogeneous collisions ($\nu=100$).
@@ IV Discussion
From the simplest one excitation case, see Sec. III.1, we found that disorder tends to suppress transport through interference, which competes with collisional noise while the latter introduces dephasing, thus facilitating delocalization, instead. One first relevant message emerging from our study is that among these competing mechanisms, low collision rates ($r_{c}\ll 1$) favor the occurrence of localized regions in the form of Inverse Participation Ratio (IPR) plateaus in time windows where collisions are absent. Interestingly, the length and rate of these plateaus depend directly on noise parameters with a universal scaling, allowing to systematically tailor sequences of transport and localization that can be applied for example to stroboscopic protocols in quantum technologies. At higher collision rates ($r_{c}\geq 1$) instead, frequent collisional events completely disrupt localization, driving the system toward a diffusive regime characterized by smooth evolution of IPR over time, even for moderate levels of disorder range.
From the case of several excitations, see Sec. III.2 and Appendix B, we learn that additional features emerge due to interactions among excitations. In particular, more localized regions in the form of plateaus appear even for the tiniest levels of disorder range.
@@ Acknowledgements.
Appendix B Appendix B. Two excitations next to each other